OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. Type your answer in the box without skipping any spaces. m = -4, point (1, -3)

OpenStudy (anonymous):

Can someone help me with this? I know it has to be in y=mx+b form

OpenStudy (anonymous):

Use point slope formula then simplify.

OpenStudy (anonymous):

y-y(initial) = m(x - x(initial))

OpenStudy (anonymous):

So, y+3=-4(x-1) simplifies to y=-4x+1

OpenStudy (anonymous):

Oh okay, thank you :)

OpenStudy (anonymous):

y+3=-4x+1

OpenStudy (anonymous):

so then y=-4x+1

OpenStudy (anonymous):

You're really pretty btw :) very beautiful. love ur hair too.

OpenStudy (anonymous):

Thank you.